# Woodstock’24: Symposium on the irreproducible science, June 07–11, 2024, Woodstock, NY

### v1

@@ Abstract
The Scholarly Hybrid Question Answering over Linked Data (QALD) Challenge at International Semantic Web Conference (ISWC) 2024 focuses on Question Answering (QA) over diverse scholarly sources: DBLP, SemOpenAlex, and Wikipedia-based texts. This paper describes a methodology that combines SPARQL queries, divide and conquer algorithms, and BERT-based-case-SQuad2 predictions. It starts with SPARQL queries to gather data, then applies divide and conquer to manage various question types and sources, and uses BERT to handle personal author questions. The approach, evaluated with Exact Match and F-score metrics, shows promise for improving QA accuracy and efficiency in scholarly contexts.
@@ 1 Introduction
The Scholarly Hybrid Question Answering over Linked Data (QALD) aims to answer questions in scholarly publications provided in natural language [6]. A challenge on Question Answering over Linked Data (QALD) which is hosted at the International Semantic Web Conference(ISWC) 2024 [5] since 2023. The 2024 edition is devoted to the development of question answering (QA) systems capable of integrating and querying information from three distinct but interconnected sources: DBLP Knowledge Graph [^1] https://dblp-april24.skynet.coypu.org/sparql, SemOpenAlex Knowledge Graph [^2] https://semoa.skynet.coypu.org/sparql, and Wikipedia-based texts [^3] https://drive.google.com/file/d/1ISxvb4q1TxcYRDWlyG-KalInSOeZqpyI/view?usp=drive_link.
@@ 2.1 Data Processing and Query Execution
The process began with executing a general script containing SPARQL queries against SemOpenAlex for both authors and institutions. This step involved querying the semoa_authors.trig file and the institution-semopenalex.trig dataset from October 2023 locally. The query execution took approximately 62-65 hours to complete due to the size and complexity of the datasets. In some cases, it filled up the RAM and returned "Killed".
Fig. 2 provides an overview of the data processing outputs from the the various KGs given. The data processing involved cleaning the dataset to remove noise such as un-added parts, mis-matched names, Names which are not well spelled, and alot more and irrelevant information such as None-useful links, . The questions were subjected to a thorough, individual examination , and keywords from the questions were noted. An example is provided by the Equation  LABEL:eq:broke-set . The following points presents the different steps of the data processing:
@@
    the datasets was transformed in alphabetical order of the questions. This allow us to see which questions had a similar structure/pattern. For questions that had a similar structure, we found that the responses were usually in the same position of the KG.
@@
    Look for the names returned by the author_dblp_uri and compare the results found in the endpoint and the results found in the Knowledge Graph (semoa_authors.trig file).
@@
$M=\\ \text{Breakdown sets}(DPR)=\left\{\begin{array}[]{l}\text{List\_author\_dblp\_uri}_{\text{set}}\\ \text{Authors}_{\text{set}}\\ \text{institution}_{\text{set}}\\ \text{hIndex}_{\text{set}}\\ \text{i10index}_{\text{set}}\\ \text{acronym}_{\text{set}}\\ \text{etc}\end{array}\right.$ (1)
@@ 2.2 Divide and Conquer Approach
    Initial Data Breakdown: We first segmented the test data based on whether the author_dblp_uri contained multiple links or a single link. This allowed us to address questions with multiple author identifiers separately from those with single identifiers.
@@
    Further Segmentation: Questions were further classified into those concerning individual authors and those about authors’ institutions. Keywords like "Organizations," "Affiliations," "institution," etc. were used to automate this classification.
@@
    Detailed Sub-Classification: Questions about authors were subdivided into more specific queries based on the information sought (e.g., publication details, affiliations).
@@ 2.4 Large Language Model-Based Predictions
The LLM used in this challenge was BERT-base-cased-squad2, a pretrained non-finetuned model download from Hugging Face [^4] https://huggingface.co/deepset/bert-base-cased-squad2, which was used for Question Answering tasks.
@@
    Integration: The LLM-generated responses were integrated with the initial query-based results before the final refinement stage to enhance the accuracy and completeness of the answers.
@@ 3 Results and Discussion
  • the BERT-base-cased-squad2 model combined with DPR algorithm significantly improved the accuracy of entity and relation extraction on the DBLP KG. It should be noted that these information are needed to provide the context for the prediction by the LLMs.
  • To handle the complete dataset, the DPR algorithm was employed, so as to be able to get through all the broken sets of the dataset.

### v2

@@ Abstract
The Scholarly Hybrid Question Answering over Linked Data (QALD) Challenge at the International Semantic Web Conference (ISWC) 2024 focuses on Question Answering (QA) over diverse scholarly sources: DBLP, SemOpenAlex, and Wikipedia-based texts. This paper describes a methodology that combines SPARQL queries, divide and conquer algorithms, and a pre-trained extractive question answering model. It starts with SPARQL queries to gather data, then applies divide and conquer to manage various question types and sources, and uses the model to handle personal author questions. The approach, evaluated with Exact Match and F-score metrics, shows promise for improving QA accuracy and efficiency in scholarly contexts.
@@ 1 Introduction
The Scholarly Hybrid Question Answering over Linked Data (QALD) aims to answer hybrid questions in scholarly publications provided in natural language [6]. The challenge focuses on Question Answering over Linked Data (QALD) and has been hosted at the International Semantic Web Conference (ISWC) 2024 [5] since 2023. The 2024 edition is devoted to the development of question answering (QA) systems capable of integrating and querying information from three distinct but interconnected sources: DBLP Knowledge Graph [^1] https://dblp-april24.skynet.coypu.org/sparql, SemOpenAlex Knowledge Graph [^2] https://semoa.skynet.coypu.org/sparql, and Wikipedia-based texts [^3] https://drive.google.com/file/d/1ISxvb4q1TxcYRDWlyG-KalInSOeZqpyI/view?usp=drive_link.
@@ 2.1 Data Processing and Query Execution
The process began with executing a general script containing SPARQL queries against SemOpenAlex for both authors and institutions. This step involved querying the semoa_authors.trig file and the institution-semopenalex.trig dataset from October 2023 locally. The query execution took approximately 62-65 hours to complete due to the size and complexity of the datasets.
Fig. 2 provides an overview of the data processing outputs from the various knowledge graphs (KGs). The data processing involved cleaning the dataset to remove noise, which included unadded parts (e.g., incomplete data fields), misspelled names, and irrelevant information such as broken or non-useful links. The knowledge graph responses were cleaned by cross-referencing returned data with authoritative sources like DBLP and SemOpenAlex to correct inconsistencies and remove redundancies. The breakdown of the cleaned data is illustrated in Equation 1, which categorizes the information into distinct sets (e.g., Authors, Institution).
An example of how the "CitedBy count" set was created from this cleaning process is shown in Fig. 3, similar to how the other broken sets (e.g., hIndex, i10index) were generated. Below are the key steps in data processing:
@@
    The datasets were transformed alphabetically according to the questions. This allowed us to identify similar structures or patterns in the questions in terms of the responses that would be returned.
@@
    We compared the names returned by the author_dblp_uri with those found in both the SPARQL endpoint results and the semoa_authors.trig knowledge graph file.
@@
$M=\\ \text{Breakdown sets}(D\&C)=\left\{\begin{array}[]{l}\text{List\_author\_dblp\_uri}_{\text{set}}\\ \text{Authors}_{\text{set}}\\ \text{institution}_{\text{set}}\\ \text{hIndex}_{\text{set}}\\ \text{i10index}_{\text{set}}\\ \text{acronym}_{\text{set}}\\ \text{etc}\end{array}\right.$ (1)
@@ 2.2 Divide and Conquer Approach
    Initial Data Breakdown: The test data were first segmented based on whether the author_dblp_uri contained multiple links or a single link. This allowed us to address questions with multiple author identifiers separately from those with single identifiers.
@@
    Further Segmentation: Questions were further classified into those concerning individual authors and those related to institutions. Keywords like "Organizations," "Affiliations," "Institution," etc. were used to automate this classification based on the content of the questions.
@@
    Detailed Sub-Classification: Within the author-related questions, further subdivisions were made depending on the specific information sought. For example, we classified questions according to whether they requested publication details, citation counts, or institutional affiliations. This sub-classification relied heavily on extracting and analyzing keywords within the questions to tailor the SPARQL queries appropriately.
    As illustrated in Equation 1, the dataset was broken into different sets such as List_author_dblp_uri, Authors, Institution, hIndex, and more. Each set represented a component that corresponded to the key information extracted from the questions. For instance, Fig. 3 provides a visual example of how the "CitedBy count" set was created. This method was applied similarly to other components, ensuring a structured approach for processing and querying different knowledge graphs.
Figure 3: Example of the "CitedBy count" set creation.
@@ 2.4 Large Language Model-Based Predictions
The LLM used in this challenge was BERT-base-cased-SQuAD2, a pretrained model fine-tuned on the Stanford Question Answering Dataset, downloaded from Hugging Face [^4] https://huggingface.co/deepset/bert-base-cased-squad2. This model was chosen without additional fine-tuning due to its capability to answer scholarly-related questions from a phrase, which aligns with our methodology for this year’s Question Answering tasks. Future work will focus on fine-tuning the LLM with this year’s dataset to evaluate its performance after fine-tuning. It should be noted that, although BERT does not have an exceptionally large number of parameters compared to more recent models, it is still considered an LLM, as described in section II of the article Large Language Models: A Survey [^5] https://arxiv.org/html/2402.06196v2.
@@
    Integration: The LLM-generated responses were integrated with the initial query-based results before the final refinement stage to enhance the accuracy and completeness of the answers. The final refinement involved combining questions resolved by the Local Predictions and those resolved using LLM-based predictions, as well as the questions found in the combined output from the Divide and Conquer(D&C) algorithms that were not present in the answers2.txt.
2.5 How questions were approached
Figure 4: Sample of how questions were approached
Fig. 4 provides an overview of the data processing outputs from the various knowledge graphs (KGs). The process of answering a question typically begins by using the author_dblp_link to retrieve the author’s name from the DBLP knowledge graph. Once the author names are identified, they undergo a series of SPARQL queries to extract relevant information about the author from both the DBLP and SemOpenAlex KGs.
For instance, the first query fetches the author’s name using their unique DBLP link, as depicted in Fig. 7. This is followed by additional queries that gather further author-specific details, such as the number of publications and citation counts, shown in Fig. 7. Finally, institution-related data about the author is retrieved, as shown in Fig. 7.
Figure 5: SPARQL query to retrieve author name
Figure 6: SPARQL query to retrieve author information
Figure 7: SPARQL query to retrieve author’s institution information
@@ 3 Results and Discussion
  • the BERT-base-cased-squad2 model combined with Divide and conquer(D&C) algorithm significantly improved the accuracy of entity and relation extraction on the DBLP KG. It should be noted that these information are needed to provide the context for the prediction by the LLMs.
  • To handle the complete dataset, the Divide and conquer(D&C) algorithm was employed, so as to be able to get through all the broken sets of the dataset.
@@ 4 Conclusion
5 Online Resources
The source code for this project is available via